OpenStudy (anonymous):

Can someone help me please??? What is the solution to x= squareroot 3x+10?

OpenStudy (anonymous):

first square both sides, becoming:\[x^2 = 3x + 10\] you can complete the square or put it in this form and use the quadratic formula:\[x^2 - 3x - 10 = 0\]

OpenStudy (anonymous):

let me know if i should keep going

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for polynomial of form: \[ax^2 + bx + c\]you can use the quadratic formula to solve for x. you will get two answers because of the plus/minus sign (one for plus one for minus) quadratic formula:\[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\] we have a = 1, b = -3 and c = -10. we simply plug them in the quadratic formula: \[x = \frac{ -(-3) \pm \sqrt{(-3)^2 - 4(1)(-10)} }{ 2(1) } = \frac{ 3 \pm \sqrt{49} }{ 2 } = \frac{ 3 \pm 7 }{ 2 }\] answers are:\[x = 5\ and\ x = -2\] let me know if you have

OpenStudy (anonymous):

thanks